OpenStudy (anonymous):

2/9 divided by 2/5 = ? MEDAL will be given.(:

OpenStudy (anonymous):

5/9 actually i think.

OpenStudy (anonymous):

Yeah 5/9

OpenStudy (jdoe0001):

\(\bf \cfrac{2}{9}\div \cfrac{2}{5}\implies \cfrac{\frac{2}{9}}{\frac{2}{5}}\implies \cfrac{2}{9}\cdot \cfrac{5}{2}\)

OpenStudy (anonymous):